# Convergence Analysis of An Alternating Nonlinear GMRES on Linear Systems

### v1

@@ Abstract
In this work, we develop an alternating nonlinear Generalized Minimum Residual (NGMRES) algorithm with depth $m$ and periodicity $p$, denoted by aNGMRES($m,p$), applied to linear systems. We provide a theoretical analysis to quantify by how much aNGMRES($m$) can improve the convergence speed of the underlying fixed-point iteration for diagonalizable and symmetric positive definite cases. Our theoretical analysis gives us a better understanding of which factors affect the convergence speed. Moreover, under certain conditions, we prove the periodic equivalence between aNGMRES applied to Richardson iteration and GMRES. Specifically, aNGMRES($\infty,p$) and full GMRES are identical at the iteration index $jp$. aNGMRES($\infty,p$) can be regarded as an alternative to GMRES for solving linear systems. For finite $m$, the iterates of aNGMRES($m,m+1$) and restarted GMRES (GMRES($m+1$)) are the same at the end of each periodic interval of length $p$, i.e, at the iteration index $jp$. The advantages of aNGMRES($m,p$) method are that there is no need to solve a least-squares problem at each iteration which can reduce the computational cost, and it can enhance the robustness against stagnations, which could occur for NGMRES($m$).
@@ 1 Introduction
where $A\in\mathbb{R}^{n\times n}$ is invertible, and $u,b\in\mathbb{R}^{n}$. Assume that $u^{*}$ is the exact solution. There are various iterative methods to solve (1), such as Jacobi iteration, Gauss-Seidel, and successive over-relaxation, Krylov subspace methods, preconditioned iterations, and multigrid methods. Acceleration methods, such as Chebyshev acceleration [4], Nesterov acceleration [10], Anderson-type acceleration [1, 21, 5, 3, 12, 2], and nonlinear GMRES (NGMRES) [22, 11], are often used to speed up the convergence of iterative methods. There is increasing interest in understanding and evaluating how acceleration methods can enhance the underlying iterative methdos and their applications. In this work, we are interested in using NGMRES to accelerate fixed-point iterations applied to (1). NGMRES was originally proposed by Washio and Oosterlee [22] to accelerate nonlinear multigrid methods. We notice that there are few studies on the development of NGMRES. In [18], the author considered steepest descent preconditioning for NGMRES, which significantly accelerates the convergence of stand-alone steepest descent optimization. While NGMRES has demonstrated competitive performance compared to other acceleration methods, its theoretical foundations, algorithmic improvements, and potential applications remain underexplored–even in the content of linear problems. Our goal is to better understand NGMRES applied to linear problems, and propose a new variant of NGMRES.
@@
First, we provide a brief review of NGMRES. NGMRES with a finite window size $m$ (i.e., using the previous $m+1$ iterates), denoted by NGMRES($m$), is preferred over full version because of its greater computational efficiency. In 1997, Washio and Oosterlee [22] proposed Krylov subspace acceleration analogous to GMRES for nonlinear multigrid on the finest level. In 2000, Oosterlee and Washio in [11] proposed a parallel nonlinear Krylov acceleration strategy for solving nonlinear equations, which utilized multigrid and GMRES methods. In [17, 19], NGMRES is applied to accelerate the alternating least squares for canonical tensor decomposition problems. However, there is little mathematical understanding of theory or solution approaches for nonlinear problems with NGMRES. Analysis of the asymptotic convergence of NGMRES can provide a characterization of behavior of NGMRES. There are fundamental gaps in our current understanding of how NGMRES works, and what the relationships between NGMRES, GMRES [15], and restarted GMRES (GMRES($m$)) [16] are. In [8], the worst-case root-convergence of GMRES(1) has been proposed. In fact, NGMRES(0) is GMRES(1). Thus, these results in [8] serve an upper bound on the performance of NGMRES($m$) on linear systems. Assume that the 2-norm of the residual of GMRES is strictly decreasing. We build connection between NGMRES and GMRES when applied to linear systems. In [6], we proved that the residuals of NGMRES($\infty$) are the same as these of GMRES, and if $A$ is invertible, the iterates generated by the two methods are identical. Moreover, for $A$ being symmetric, we proved that NGMRES($m$) and GMRES are identical for $\forall m>1$, indicating no need to use large window size $m$. Specially, when $A$ is symmetric, NGMRES(1) is equivalent to the conjugate residual method. These discoveries help us better understand NGMRES on linear problems. For nonlinear problems, we have provided convergence analysis for NGMRES($m$) provided that the underlying fixed-point operator is contractive [7]. Specifically, we proved that for general $m>0$, the residuals of NGMRES($m$) converge r-linearly under some conditions. For $m=0$, we proved that the residuals of NGMRES(0) converge q-linearly. In practice, we see NGMRES dramatically enhances the performance of linear and nonlinear fixed-point iterations. However, there is a lack of studies that quantitatively assess the improvements achieved by NGMRES compared to the underlying fixed-point iteration. One aim of this work is to quantify the improvement achieved by considering a single step of NGMRES.
@@
In this work, we first study how NGMRES can improve the performance of a fixed-point iteration by considering one-step NGMRES, that is, given $u_{0}$, we compute $u_{j}=q(u_{j-1})$, where $j=1,2,\cdots,k$, and then generate $u_{k+1}$ by NGMRES($m$). Note that at each iteration of NGMRES($m$), one needs to solve a least-squares problem, which might be computational expensive for large $m$. Then, based on one-step analysis of NGMRES, we propose alternating NGMRES with depth $m$ and periodicity $p$, denoted by aNGMRES($m,p$). The main idea is that, given an initial guess $x_{0}$, we compute $p-1$ fixed point iterations, then switch to a single NGMRES($m$) iteration. This process is then repeated. We provide convergence analysis of aNGMRES, and make connections between aNGMRES, GMRES, and restarted GMRES. Additionally, we make comments on recently proposed alternating Anderson-Richardson method (AAR). Our proposed aNGMRES method offers a novel approach to solving large linear systems. It can be extended to accelerate nonlinear fixed-point iterations.
@@
The remainder of this work is organized as follows. In section 2, we introduce NGMRES and analyze the improvement of one step NGMRES. In section 3, we propose aNGMRES, establish relationships with full GMRES and restarted GMRES, and provide convergence anlaysis of aNGMRES. In section 4, we present some numerical examples to validate our theoretical findings. We draw conclusions in section 5.
@@ 2 NGMRES
NGMRES [22, 11] was originally developed for acceleration nonlinear iterations. Here, we consider NGMRES applied to linear systems. Assume we have the following fixed-point iteration for (1)
@@
where $M=I-A$. Note that (2) is the Richardson iteration. Define the $k$th residual as
$r_{k}=r(u_{k})=u_{k}-q(u_{k})=Au_{k}-b.$
[…]
@@
    $u_{k+1}=q(u_{k})+\sum_{i=0}^{m_{k}}\beta^{(k)}_{i}\left(q(u_{k})-u_{k-i})\right),$
    where $\beta_{i}^{(k)}$ is obtained by solving the following least-squares problem
    $\min_{\beta_{i}^{(k)}}\|r(q(u_{k}))+\sum_{i=0}^{m_{k}}\beta^{(k)}_{i}\left(r(q(u_{k}))-r(u_{k-i})\right)\|^{2}.$ (3)
@@
From [6], it reveals that the least-squares problem (3) is minimizing the $(k+1)$th residual. We state the result in the following.
@@
$\|r_{k+1}\|=\|Au_{k+1}-b\|=\min_{\beta_{i}^{(k)}}\|r(q(u_{k}))+\sum_{i=0}^{m_{k}}\beta^{(k)}_{i}\left(r(q(u_{k}))-r(u_{k-i})\right)\|^{2},$ (4)
@@
$\|r_{k+1}\|\leq\|r_{k}\|.$ (5)
Since the 2-norm of the residuals of NGMRES($m$) is nonincreasing, NGMRES($m$) always converges. Let $\gamma_{j}^{(k)}=\sum_{i=j}^{m_{k}}\beta_{i}^{(k)}$. For the convenience of convergence analysis, we rewrite Algorithm 1 as Algorithm 2.
@@
    $\min_{\gamma_{i}^{(k)}}\|r(q(u_{k}))+\gamma_{0}^{(k)}(r(q(u_{k}))-r(u_{k}))+\sum_{i=1}^{m_{k}}\gamma^{(k)}_{i}\left(r(u_{k-i+1})-r(u_{k-i})\right)\|^{2}.$ (7)
@@
Let
$\bm{\gamma}^{(k)}=\begin{bmatrix}\gamma_{m_{k}}^{(k)},&\gamma_{m_{k}-1}^{(k)},&\cdots&\gamma_{0}^{(k)}\end{bmatrix}^{T}.$
and define the following matrices with dimensions $n\times m_{k}$,
$\mathcal{S}_{k}=\begin{bmatrix}u_{k-m_{k}+1}-u_{k-m_{k}},&u_{k-m_{k}+2}-u_{k-m_{k}+1},&\cdots&u_{k}-u_{k-1},&q(u_{k})-u_{k}\end{bmatrix}.$
@@
$\min_{\bm{\gamma}^{(k)}}\|r(q(u_{k}))+\mathcal{D}_{k}\bm{\gamma}^{(k)}\|^{2}_{2},$
@@
$\bm{\gamma}^{(k)}=-(\mathcal{D}_{k}^{T}\mathcal{D}_{k})^{-1}\mathcal{D}_{k}^{T}r(q(u_{k})).$
@@
$\mathcal{S}_{k}=\begin{bmatrix}e_{k-m_{k}+1}-e_{k-m_{k}}&e_{k-m_{k}+2}-e_{k-m_{k}+1},&\cdots,&e_{k}-e_{k-1},&-Ae_{k}\end{bmatrix}.$
@@
$\mathcal{D}_{k}=A\begin{bmatrix}e_{k-m_{k}+1}-e_{k-m_{k}}&e_{k-m_{k}+2}-e_{k-m_{k}+1},&\cdots,&e_{k}-e_{k-1},&-Ae_{k}\end{bmatrix}=AS_{k}.$
@@ 2.1 Improvement of NGMRES
Here, we study how NGMRES($m$) can improve the performance of a fixed-point iteration. In order to do this, we consider one-step NGMRES, that is, given $u_{0}$, we compute $u_{j}=q(u_{j-1})$, where $j=1,2,\cdots,k$, and then generate $u_{k+1}$ by NGMRES($m$). We compare $u_{k+1}$ with $q(u_{k})$ which is the fixed-point iterate.
@@
[15] Let $[\alpha,\beta]$ be a non-empty interval in $\mathbb{R}$, which does not include 00 and 1, and let $p_{s}(t)$ be a polynomial with degree at most $s$. Then,
$\min_{p_{s}\in P_{s},p_{s}(\gamma)=1}\max_{t\in[\alpha,\beta]}|p_{s}(t)|=\frac{1}{|C_{s}(1+2\frac{\gamma-\beta}{\beta-\alpha})|},$
@@
Let $a<b,\alpha=\frac{1}{b}$ and $\beta=\frac{1}{a}$ and $[\alpha,\beta]$ does not include 00 and 1. Denote
$\epsilon(a,b,s)=\min_{p_{s}\in P_{s},p_{s}(1)=1}\max_{t\in[\alpha,\beta]}|p_{s}(t)|=\frac{1}{|C_{s}(1+2\frac{1-\beta}{\beta-\alpha})|}.$
@@ 2.1.1 Diagonalizable case
$=A^{-1}(I-UH(H^{T}U^{T}UH)^{-1}H^{T}U^{T})A$
$=U(I-\Lambda)^{-1}U^{-1}(I-UH(H^{T}U^{T}UH)^{-1}H^{T}U^{T})U(I-\Lambda)U^{-1}$
@@
$G_{k}=\Theta(I-\mathcal{H}(\mathcal{H}^{T}U^{T}U\mathcal{H})^{-1}\mathcal{H}^{T}U^{T}U)\Theta^{-1},\quad\Theta=(I-\Lambda)^{-1}$ (11)
@@
Assume that $M$ is a diagonalizable matrix, i.e., $M=U\Lambda U^{-1}$, where $\Lambda$ is a diagonal matrix, the eigenvalues of $M$ are contained in an interval $[a,b]$ that does not include 00 and 1, and $m\leq k$. Then, the 2-norm of the residual $r_{k+1}$ obtained from one-step NGMRES($m$) satisfies
$\|r_{k+1}\|\leq\epsilon(a,b,m+1)\cdot\kappa_{2}(U)\cdot\|Mr_{k}\|.$ (12)
where $\kappa_{2}(U)=\|U\|_{2}\|U^{-1}\|$.
Note that if the $(k+1)$th iterate is obtained by fixed-point iteration, then the corresponding residual is $Mr_{k}$. Our theoretical result provides a quantitative measurement of how much the fixed-point iteration can be improved. In fact, the factor $\epsilon(a,b,m+1)\cdot\kappa_{2}(U)$ is the gain in how much NGMRES($m$) improves the convergence speed of the fixed-point iteration.
@@
$\leq\|U\|\|U^{-1}\|\min_{p_{m}\in P_{m}}\|I-p_{m}(\Lambda)(I-\Lambda)\Lambda^{-m-1}\|_{2}\|Mr_{k}\|$
$\leq\|U\|\|U^{-1}\|\min_{p_{m}\in P_{m}}\max_{t\in[a,b]}\|I-p_{m}(t)t^{-m-1}(1-t)\|_{2}\|Mr_{k}\|$
$\leq\|U\|\|U^{-1}\|\min_{p_{m}\in P_{m}}\max_{t\in[1/b,1/a]}\|I-p_{m}(1/t)t^{m}(t-1)\|\|Mr_{k}\|$
$\leq\|U\|\|U^{-1}\|\min_{p_{m}\in P_{m}}\max_{t\in[1/b,1/a]}\|I-p_{m}(t)(t-1)\|_{2}\|Mr_{k}\|$
$\leq\|U\|\|U^{-1}\|\min_{q_{m+1}\in P_{m+1},q_{m+1}(1)=1}\max_{t\in[1/b,1/a]}|q_{m+1}(t)|\|Mr_{k}\|.$
@@
We note that the estimate (13) is the same as the residual norm of the $(m+1)$th step of GMRES, see [15]. In section 3, we will see that under certain condition, the $(m+1)$th iteration, $u_{m+1}$, generated by aNGMRES($\infty,m+1$)–the method proposed in this work–is identical to $(m+1)$th iteration of GMRES.

### v2

@@ Abstract
In this work, we develop an alternating nonlinear Generalized Minimum Residual (NGMRES) algorithm with depth $m$ and periodicity $p$, denoted by aNGMRES($m,p$), applied to linear systems. We provide a theoretical analysis to quantify by how much one-step NGMRES($m$) using Richardson iterations as initial guesses can improve the convergence speed of the underlying fixed-point iteration for diagonalizable and symmetric positive definite cases. Our theoretical analysis gives us a better understanding of which factors affect the convergence speed. Moreover, under certain conditions, we prove the periodic equivalence between the proposed aNGMRES applied to Richardson iteration and GMRES. Specifically, aNGMRES($\infty,p$) and full GMRES are identical at the iteration index $jp$. Therefore, aNGMRES($\infty,p$) can be regarded as an alternative to GMRES for solving linear systems. For finite $m$, the iterates of aNGMRES($m,m+1$) and restarted GMRES (GMRES($m+1$)) are the same at the end of each periodic interval of length $p$, i.e, at the iteration index $jp$. In Addition, we present a convergence analysis of aNGMRES when applied to accelerate Richardson iteration. The advantages of aNGMRES($m,p$) method are that there is no need to solve a least-squares problem at each iteration which can reduce the computational cost, and it can enhance the robustness against stagnations, which could occur for NGMRES($m$).
@@ 1 Introduction
where $A\in\mathbb{R}^{n\times n}$ is invertible, and $u,b\in\mathbb{R}^{n}$. Assume that $u^{*}$ is the exact solution. There are various iterative methods to solve (1), such as Jacobi iteration, Gauss-Seidel, successive over-relaxation, Krylov subspace methods, preconditioned iterations, and multigrid methods. Acceleration methods, such as Chebyshev acceleration [7], Nesterov acceleration [16], Anderson-type acceleration [1, 24, 8, 3, 18, 2], and nonlinear GMRES (NGMRES) [25, 17], are often used to speed up the convergence of iterative methods. There is increasing interest in understanding and evaluating how acceleration methods can enhance the underlying iterative methods and their applications. In this work, we are interested in using NGMRES to accelerate fixed-point iterations applied to (1), and in analyzing its convergence properties as well as its connections to existing methods. NGMRES was originally proposed by Washio and Oosterlee [25] to accelerate nonlinear multigrid methods. We note that there are few studies on the development of NGMRES. In De Sterck’s work [5], steepest descent preconditioning for NGMRES was examined, which significantly accelerates the convergence of stand-alone steepest descent optimization. While NGMRES has demonstrated competitive performance compared to other acceleration methods, its theoretical foundations, algorithmic improvements, and potential applications remain underexplored—even in the content of linear problems. Our goal is to better understand NGMRES applied to linear systems, and to propose a new variant of NGMRES.
@@
First, we provide a brief review of NGMRES. NGMRES with a finite window size $m$ (i.e., using the previous $m+1$ iterates), denoted by NGMRES($m$), is preferred over full version because of its greater computational efficiency. In 1997, Washio and Oosterlee [25] proposed Krylov subspace acceleration analogous to GMRES for the nonlinear multigrid on the finest level. In 2000, Oosterlee and Washio [17] proposed a parallel nonlinear Krylov acceleration strategy for solving nonlinear equations, which utilized multigrid and GMRES methods. Later, NGMRES [4, 6] is applied to accelerate the alternating least squares for canonical tensor decomposition problems. However, there is little mathematical understanding of the theory or solution approaches for nonlinear problems with NGMRES. Analysis of the asymptotic convergence of NGMRES can provide a characterization of the behavior of NGMRES. There are fundamental gaps in our current understanding of how NGMRES works, and what the relationships between NGMRES, GMRES [21], and restarted GMRES (GMRES($m$)) [22] are. The worst-case root-convergence of GMRES(1) has been proposed [12]. In fact, NGMRES(0) is GMRES(1). Thus, the results of GMRES(1) [12] serve as an upper bound on the performance of NGMRES($m$) on linear systems. Recently, we built a connection between NGMRES and GMRES when applied to linear systems. Assume that the 2-norm of the residual of GMRES is strictly decreasing. We [10] proved that the residuals of NGMRES($\infty$) are the same as these of GMRES, and if $A$ is invertible, the iterates generated by the two methods are identical. Moreover, for invertible matrices $A$ that are either symmetric or shifted skew-symmetric of the form $A=\alpha I+S$, where $\alpha\in\mathbb{R}$ and $S$ is skew-symmetric, we proved that NGMRES($m$) and GMRES are identical for $\forall m\geq 1$, suggesting that a large window size $m$ is unnecessary. Specially, when $A$ is symmetric and invertible, NGMRES(1) is equivalent to the conjugate residual method. These discoveries help us better understand NGMRES on linear problems. For nonlinear problems, we have provided convergence analysis for NGMRES($m$) for a specific class of contractive fixed-point operators[11]. Specifically, we proved that for general $m>0$, the residuals of NGMRES($m$) converge r-linearly under some conditions. For $m=0$, we proved that the residuals of NGMRES(0) converge q-linearly. In practice, we see NGMRES dramatically enhances the performance of linear and nonlinear fixed-point iterations. However, there is a lack of studies that quantitatively assess the improvements achieved by NGMRES compared to the underlying fixed-point iteration. One aim of this work is to quantify the improvement achieved by considering a single step of NGMRES applied to linear systems using Richardson iterations.
@@
In this work, we first study how NGMRES can improve the performance of a fixed-point iteration by considering one-step NGMRES on linear problems, that is, given $u_{0}$, we compute $u_{j}=q(u_{j-1})$, where $j=1,2,\cdots,k$, and then generate $u_{k+1}$ by NGMRES($m$). Note that at each iteration of NGMRES($m$), one needs to solve a least-squares problem, which might be computational expensive for large $m$. Then, based on one-step analysis of NGMRES, we propose alternating NGMRES with depth $m$ and periodicity $p$, denoted by aNGMRES($m,p$). The main idea is that, given an initial guess $x_{0}$, we compute p1p-1 fixed point iterations, then switch to a single NGMRES($m$) iteration. This process is then repeated. We provide convergence analysis of aNGMRES, and make connections between aNGMRES, GMRES, and restarted GMRES. Additionally, we make comments on recently proposed alternating Anderson-Richardson method (AAR). Our proposed aNGMRES method offers a novel approach to solving large linear systems. It can be extended to accelerate nonlinear fixed-point iterations.
@@
The remainder of this work is organized as follows. In section 2, we introduce NGMRES and analyze the improvement of one- step NGMRES in two distinct cases: (1) M=IAM=I-A is diagonalizable, and (2) when $M$ is symmetric and positive definite. In section 3, we propose aNGMRES, establish relationships with full GMRES and restarted GMRES, and provide convergence analysis of aNGMRES. In section 4, we present some numerical examples to validate our theoretical findings. Additionally, we discuss the practical considerations of aNGMRES and present the performance of aNGMRES applied to both linear and nonlinear problem settings. We draw conclusions in section 5.
@@ 2 NGMRES
NGMRES [25, 17] was originally developed to accelerate nonlinear iterations. Here, we consider NGMRES applied to linear systems. Assume we have the following fixed-point iteration for (1)
@@
where M=IAM=I-A. Note that (2) is the Richardson iteration. Define the $k$th residual as
$r_{k}=r(u_{k})=u_{k}-q(u_{k})=Au_{k}-b.$ (3)
In practice, the fixed-point iteration (2) converges slowly or even diverges. We consider NGMRES($m$) to accelerate (2). In each iteration, a least-squares problem must be solved to determine the combination coefficients. We present the framework of NGMRES($m$) in Algorithm 1. Throughout this work, $\|\cdot\|$ denotes the 2-norm ($\ell^{2}$) of a vector or a matrix.
@@
    $u_{k+1}=q(u_{k})+\sum_{i=0}^{m_{k}}\beta^{(k)}_{i}\left(q(u_{k})-u_{k-i})\right),$ (4)
    where ${\bm{\beta}^{(k)}}=[\beta_{0}^{(k)},\beta_{1}^{(k)},\cdots,\beta_{m_{k}}^{(k)}]^{T}$ is obtained by solving the following least-squares problem
    $\min_{\bm{\beta}^{(k)}}\|r(q(u_{k}))+\sum_{i=0}^{m_{k}}\beta^{(k)}_{i}\left(r(q(u_{k}))-r(u_{k-i})\right)\|^{2}.$ (5)
@@
Remark 2.1.
We note that, when applying NGMRES to solve the general equation $g(u)=0$, the corresponding least-squares problem in Algorithm 1 takes the form[10]
$\min_{\bm{\beta}^{(k)}}\|g(q(u_{k}))+\sum_{i=0}^{m_{k}}\beta^{(k)}_{i}\left(g(q(u_{k}))-g(u_{k-i})\right)\|^{2}.$
In this work, we focus on solving $Au=b$, which can be reformulated as g(u)=Aub=0g(u)=Au-b=0, using the fixed-point iteration defined in (2). From (3), we observe that $r(u_{k})=g(u_{k})$, which justifies our use of the least-squares formulation in (5) throughout this work.
The work of Greif and He[10] shows that the least-squares problem (5) is minimizing the $(k+1)$th residual. For completeness, we restate their key results as a lemma below.
@@
$\|r_{k+1}\|=\|Au_{k+1}-b\|=\min_{\bm{\beta}^{(k)}}\|r(q(u_{k}))+\sum_{i=0}^{m_{k}}\beta^{(k)}_{i}\left(r(q(u_{k}))-r(u_{k-i})\right)\|^{2},$ (6)
@@
$\|r_{k+1}\|\leq\|r_{k}\|.$
We omit the proof and refer the reader to the work of Greif and He[10] for further details. Since the 2-norm of the residuals of NGMRES($m$) is nonincreasing, NGMRES($m$) always converges. Let $\gamma_{j}^{(k)}=\sum_{i=j}^{m_{k}}\beta_{i}^{(k)}$ and $\bm{\gamma}^{(k)}=[\gamma_{m_{k}}^{(k)},\gamma_{m_{k}-1}^{(k)},\cdots,\gamma_{0}^{(k)}]^{T}$. For the convenience of convergence analysis, we rewrite Algorithm 1 as Algorithm 2.
@@
    $\min_{\bm{\gamma}^{(k)}}\|r(q(u_{k}))+\gamma_{0}^{(k)}(r(q(u_{k}))-r(u_{k}))+\sum_{i=1}^{m_{k}}\gamma^{(k)}_{i}\left(r(u_{k-i+1})-r(u_{k-i})\right)\|^{2}.$ (8)
@@
To carry out the convergence analysis of NGMRES, our first objective is to establish the relationship between consecutive residuals or errors. Now we define the following matrices with dimensions $n\times(m_{k}+1)$,
$\mathcal{S}_{k}=\begin{bmatrix}u_{k-m_{k}+1}-u_{k-m_{k}},&u_{k-m_{k}+2}-u_{k-m_{k}+1},&\cdots&u_{k}-u_{k-1},&q(u_{k})-u_{k}\end{bmatrix}$,
@@
$\min_{\bm{\gamma}^{(k)}}\|r(q(u_{k}))+\mathcal{D}_{k}\bm{\gamma}^{(k)}\|^{2}_{2},$ (9)
@@
$\bm{\gamma}^{(k)}=-(\mathcal{D}_{k}^{T}\mathcal{D}_{k})^{-1}\mathcal{D}_{k}^{T}r(q(u_{k})).$ (10)
@@
$\mathcal{S}_{k}=\begin{bmatrix}e_{k-m_{k}+1}-e_{k-m_{k}}&e_{k-m_{k}+2}-e_{k-m_{k}+1},&\cdots,&e_{k}-e_{k-1},&-Ae_{k}\end{bmatrix}$,
@@
$\mathcal{D}_{k}=A\begin{bmatrix}e_{k-m_{k}+1}-e_{k-m_{k}}&e_{k-m_{k}+2}-e_{k-m_{k}+1},&\cdots,&e_{k}-e_{k-1},&-Ae_{k}\end{bmatrix}=A\mathcal{S}_{k}.$
@@ 2.1 Improvement of NGMRES
Here, we study how NGMRES($m$) can improve the performance of the fixed-point iteration given by (2). In order to do this, we consider one-step NGMRES, that is, given $u_{0}$, we compute $u_{j}=q(u_{j-1})$, where $j=1,2,\cdots,k$, and then generate $u_{k+1}$ by NGMRES($m$). We compare $u_{k+1}$ with $q(u_{k})$ which is the fixed-point iterate.
@@
[21] Let $\alpha,\beta]$ be a non-empty interval in $\mathbb{R}$, which does not include 0 and 1, and let $p_{s}(t)$ be a polynomial with degree at most $s$. Then,
$\min_{p_{s}\in P_{s},p_{s}(\gamma)=1}\max_{t\in[\alpha,\beta]}|p_{s}(t)|=\frac{1}{\big|C_{s}(1+2\frac{\gamma-\beta}{\beta-\alpha})\big|},$
@@
Let $a_{1}<a_{2},\alpha=\frac{1}{a_{2}}$ and $\beta=\frac{1}{a_{1}}$ and $\alpha,\beta]$ does not include 0 and 1. Denote
$\epsilon(a_{1},a_{2},s)=\min_{p_{s}\in P_{s},p_{s}(1)=1}\max_{t\in[\alpha,\beta]}|p_{s}(t)|=\frac{1}{\big|C_{s}(1+2\frac{1-\beta}{\beta-\alpha})\big|}.$
We begin by examining the improvement under the assumption that $M$ is diagonalizable, followed by an analysis of the case where $M$ is symmetric and positive definite.
@@ 2.1.1 Diagonalizable case
$=A^{-1}(I-U\mathcal{H}(\mathcal{H}^{T}U^{T}U\mathcal{H})^{-1}\mathcal{H}^{T}U^{T})A$
$=U(I-\Lambda)^{-1}U^{-1}(I-U\mathcal{H}(\mathcal{H}^{T}U^{T}U\mathcal{H})^{-1}\mathcal{H}^{T}U^{T})U(I-\Lambda)U^{-1}$
@@
$G_{k}=\Theta(I-\mathcal{H}(\mathcal{H}^{T}U^{T}U\mathcal{H})^{-1}\mathcal{H}^{T}U^{T}U)\Theta^{-1},\quad\Theta=(I-\Lambda)^{-1}.$ (14)
@@
Based on the above analysis, we next present an error estimate for the residual of one-step NGMRES($m$) under the assumption that the eigenvalues of $M$ lie within a specified interval.
@@
Assume that $M$ is a diagonalizable matrix, i.e., $M=U\Lambda U^{-1}$, where $\Lambda$ is a diagonal matrix, the eigenvalues of $M$ are contained in an interval $a_{1},a_{2}]$ that does not include 0 and 1, and $m\leq k$. Then, the 2-norm of the residual $r_{k+1}$ obtained from one-step NGMRES($m$) satisfies
$\|r_{k+1}\|\leq\epsilon(a_{1},a_{2},m+1)\cdot\kappa_{2}(U)\cdot\|Mr_{k}\|.$ (15)
where $\kappa_{2}(U)=\|U\|\|U^{-1}\|$.
Note that if the $(k+1)$th iterate is obtained by the fixed-point iteration, then the corresponding residual is $Mr_{k}$. Our theoretical result (15) provides a quantitative measurement of how much the fixed-point iteration can be improved. In fact, the factor $\epsilon(a_{1},a_{2},m+1)\cdot\kappa_{2}(U)$ is the gain in how much one-step NGMRES($m$) improves the convergence speed of the fixed-point iteration.
@@
$\leq\|U\|\|U^{-1}\|\min_{p_{m}\in P_{m}}\|I-p_{m}(\Lambda)(I-\Lambda)\Lambda^{-m-1}\|\|Mr_{k}\|$
$\leq\|U\|\|U^{-1}\|\min_{p_{m}\in P_{m}}\max_{t\in[a_{1},a_{2}]}\left|1-p_{m}(t)t^{-m-1}(1-t)\right|\|Mr_{k}\|$
$\leq\|U\|\|U^{-1}\|\min_{p_{m}\in P_{m}}\max_{t\in[1/a_{2},1/a_{1}]}\left|1-p_{m}(1/t)t^{m}(t-1)\right|\|Mr_{k}\|$
$\leq\|U\|\|U^{-1}\|\min_{p_{m}\in P_{m}}\max_{t\in[1/a_{2},1/a_{1}]}\left|1-p_{m}(t)(t-1)\right|\|Mr_{k}\|$
$\leq\|U\|\|U^{-1}\|\min_{q_{m+1}\in P_{m+1},q_{m+1}(1)=1}\max_{t\in[1/a_{2},1/a_{1}]}\left|q_{m+1}(t)\right|\|Mr_{k}\|.$
@@
Next, we present an error estimate for the residual of one-step NGMRES($m$), assuming the eigenvalues of $M$ lie within a generic domain in the complex plane.
@@
We note that the estimate (16) is the same as the residual norm of the $(m+1)$th step of GMRES; see Proposition 6.32 in Saad’s work [21] for further details. In Addition, Saad’s work [21] presents the convergence behavior of GMRES when the eigenvalues of AA lie within an ellipse, which may offer useful insights for estimating (17). However, we do not pursue a detailed estimation of (17). In section 3, we will see that under certain conditions, the $(m+1)$th iteration, $u_{m+1}$, generated by aNGMRES($\infty,m+1$)—the method proposed in this work—is identical to the $(m+1)$th iteration of GMRES. In the next section, we discuss the estimation of errors and residuals for a special case of $M$.